OpenStudy (anonymous):

What are the factors of x^2 - 49? (x - 1)(x + 49) (x - 7)(x - 7) (x - 7)(x + 7) Prime This one is C right?

OpenStudy (anonymous):

Alright so this is a difference of squares. So basically what is the square root of 49?

OpenStudy (anonymous):

Difference of squares will always go like this \[a ^{2}-b ^{2} = (a-b)(a+b)\]

OpenStudy (anonymous):

Since you know this is a difference of squares (both the x^2 and the 49 are perfect squares, the x^2 is a perfect square of x, and 49 is a perfect square of 7) This means it is (x-7)(x+7)

OpenStudy (anonymous):

For example if you had x^2-36, this is really quickly (x-6)(x+6)

OpenStudy (anonymous):

Or x^2-16 (x-4)(x+4)

OpenStudy (anonymous):

Thanks. Sorry i walked away from my computer for a minute

OpenStudy (anonymous):

So any time you have a problem where it is x^2-(A Perfect Square), it goes like this Your perfect squares are 1^2, 2^2, 3^2, 4^2, 5^2, 6^2 etc So 1^1 = 1 2^2 = 4 3^2 = 9 4^2 = 16 etc.. So when you have x^2 - 1 - > (x-1)(x+1) x^2 - 4 -> (x-2)(x+2) x^2 - 9 -> (x-3)(x+3) x^2 - 16 -> What is this one?

OpenStudy (anonymous):

Select the factors of x^2 - 14x + 49. (x + 49)(x + 1) (x + 7)(x + 7) (x - 7)(x - 7) (x - 49)(x - 1) Isnt this one C?

OpenStudy (anonymous):

Read what I wrote

OpenStudy (anonymous):

This one is a perfect square trinomial. Not a difference of squares. See how its a minus in the middle and a + at the end? This means we will have 2 (-)'s. So the square here is 7. And since we need two (-)'s it is (x-7)(x-7) =0)

OpenStudy (anonymous):

It would be (x-4)(x+4) right? For the one you just asked me?

OpenStudy (anonymous):

The answer to my question above, that's exactly right!

OpenStudy (anonymous):

So what is x^2-81?

OpenStudy (anonymous):

Quick!

OpenStudy (anonymous):

What is 81 the square of?

OpenStudy (anonymous):

(x-9)(x+9)

OpenStudy (anonymous):

You freakin' got it.

OpenStudy (anonymous):

Nice

OpenStudy (anonymous):

Brb, coffee and bathoom lol

OpenStudy (anonymous):

Graph of a cubic polynomial that falls to the left and rises to the right with x intercepts negative 3, negative 2, and 2 Which of the following functions best represents the graph? f(x) = x^3 + x^2 - 4x - 4 f(x) = x^3 + x^2 - x - 1 f(x) = x^3 + 3x^2 - 4x - 12 f(x) = x^3 + 2x^2 - 6x - 12

OpenStudy (anonymous):

Okay no problem. Lol

OpenStudy (anonymous):

To find this one: \((x-@)(x-@)(x-@)\) Where @ are your x intercepts, which in this case are -3,-2, and 2 (x+3)(x+2)(x-2) Multiply all these to get your answer

OpenStudy (anonymous):

so x^2-12?

OpenStudy (anonymous):

First you can multiply the first 2 using foil: